OpenStudy (anonymous):

The initial value of a car is 35000. After a year, the value of a car is 31850. Write an exponential function to model the expected value of the car. Estimate the value of the car after 5 years.. I got 35005.49 is that correct? If not please help me!

OpenStudy (anonymous):

\[35000\times \left(\frac{31850}{35000}\right)^t\] should do it

OpenStudy (anonymous):

probably easier to write this as \[35000\times(.91)^t\]

OpenStudy (anonymous):

then \(35000\times (.91)^5\) and a calculator

OpenStudy (anonymous):

the answer you wrote cannot be correct because your car has gained value over 5 years, but it is losing value